# Who Bears the Cost? High-Frequency Evidence on Minimum Wage Effects and Amenity Pass-Through in Spot Labor Markets

### v1

@@ Abstract
This paper provides causal evidence on the short-run effects of minimum wage hikes in a spot labor market using confidential contract-level data from a Japanese job-matching platform. Leveraging high-frequency variation and a bin-level difference-in-differences design, we find a 2 percentage point decline in employment within affected wage bins, with losses concentrated below the new threshold. Transportation reimbursement—a flexible nonwage amenity—remains largely unchanged, indicating limited pass-through. These findings suggest that spot labor markets adjust rapidly through wage compliance, with the cost burden borne by establishments rather than offset through reductions in job amenities.
@@ 1 Introduction
In this paper, we use confidential contract-level data from a Japanese spot labor market platform to provide causal evidence on the effects of minimum wage hikes on short-term employment and transportation reimbursement. The former captures the immediate response of firms to minimum wage increases, while the latter, as an easily adjustable nonwage amenity, enables us to examine potential pass-through effects of minimum wage changes on broader compensation structures beyond base wages. Understanding these two outcomes is essential for capturing the impact of minimum wage hikes, as they reflect both the direct and indirect margins through which firms adjust labor conditions in response to wage floor policies.
@@
Our findings reveal that the impact of the minimum wage increase on employment in the lower end of the wage distribution. We find that the increase resulted in approximately 2 percentage point decline in total employment among wage bins directly affected by the policy, with a 3 percentage point reduction in jobs below the new minimum wage and a 1 percentage point increase just above it. These effects materialize immediately after the policy takes effect and remain stable throughout the post-reform period. Unlike evidence from broader labor markets where excess jobs offset missing jobs (Cengiz et al. 2019), our setting shows that missing jobs consistently exceed excess jobs, resulting in a net negative employment effect. These results support the view that the spot labor market adjusts quickly to minimum wage hikes, primarily through employment reallocation.
@@ 1.1 Related Literature
The rest of the paper is structured as follows: Section 2 outlines the context of the spot labor market and the Japanese Minimum Wage Act with related studies. Section 3 introduces our data and illustrates motivating data patterns. Section 4 presents our identification strategy. Section 5 reports our empirical findings. Section 6 concludes the paper.
@@ 2.2 Spot work platform
We use a term “spot work platform” to refer non-regular labor market provided by the huge online platform in Japan. [^7] According to the Japan Spot Work Association (JASWA), “spot work” refers to employment contracts for short-hour, one-off work engagements that are facilitated through digital platforms. These platforms serve as intermediaries that match job offers with available workers—often on a same-day or near-term basis—and rely on digital technologies to streamline the processes of matching, contracting, and compensation. JASWA plays a central role in promoting sound practices in the sector by issuing guidelines for legal compliance, offering interpretations of labor laws, and supporting proper labor management. In particular, it emphasizes the use of digital infrastructure to reduce the administrative burden traditionally associated with employment, including time tracking, wage calculation, and performance evaluation. This description draws on information provided by the association’s official overview and published guidelines (https://www.jaswa.or.jp/about/; https://jaswa.or.jp/pdf/spotwork_gudeline_v2.pdf). This platform operates as an on-demand staffing platform designed to connect businesses with temporary workers for short-term jobs through the app. The primary users of the platform are individuals seeking flexible, task-based employment rather than long-term positions and establishments seeking for short-term workers. Workers can register on the platform for free and immediately gain access to job offers from various companies, with the platform streamlining the process of matching workers with available shifts across industries such as food, retail, and logistics.
@@ 4.1 Prefecture-month-wage-bin level DID estimate
$\frac{Y_{j(p),t}}{N_{j(p),t}}=\sum_{l\neq-1}\sum_{e\neq\infty}\mu_{l,e}\mathbf{1}\{t-7=l\}\mathbf{1}\{E_{j(p)}=e\}+\alpha_{e}+\lambda_{t}+\epsilon_{j,t},$ (1)
@@
$\mu_{l,e}=E[Y_{j(p),7+l}-Y_{j(p),6}|E_{j(p)}=e]-E[Y_{j(p),7+l}-Y_{j(p),6}|E_{j(p)}=\infty],$
@@ 4.2 Excess-jobs-Missing-jobs Approach
Next, following Cengiz et al. (2019), we estimate the effect of minimum wage increase on the outcome throughout the wage distribution by aggregating each $\mu_{l,e}$, given by the prefecture-month-wage-bin level DID estimates. Let $\Delta a_{l}=\sum_{e=0}^{3}\mu_{l,e}$ and $\Delta b_{l}=\mu_{l,-1}$ denote the excess jobs and the missing jobs at a give relative period $l$, respectively. Let $\Delta a=\frac{1}{6}\sum_{l=0}^{5}\Delta a_{l}$ and $\Delta b=\frac{1}{6}\sum_{l=0}^{5}\Delta b_{l}$ denote the total excess jobs and missing jobs after the increase of minimum wage, respectively. We compute the overall employment effect at a give relative period $l$, denoted as $\Delta e_{l}$, by summing the excess and missing jobs at that period: $\Delta e_{l}=\Delta a_{l}+\Delta b_{l}$. Furthermore, we calculate the total employment effect after the minimum wage increase, denoted as $\Delta e$, by summing $\Delta e_{\l}$ over all relevant periods: $\Delta e=\sum_{l=0}^{5}e_{l}$.
We also estimate the employment elasticity with respect to the minimum wage as follows:
%$\frac{\%\Delta\text{Total Employment}}{\%\Delta MW}=\frac{\Delta a+\Delta b}{\%\Delta MW},$ (2)
where %$\%\Delta MW$ represents the percent change in the minimum wage averaged across prefectures.
@@ 5.1 Prefecture-month-wage-bin Level DID Estimate on Employment
Three key findings emerge from Panel (a) and Panel (b). First, immediately after the minimum wage increase, there is a clear occurrence of job losses (missing jobs) below the minimum wage and job creation (excess jobs) above the minimum wage in the wage distribution. Second, the estimated excess jobs and missing jobs remain stable across months following the minimum wage increase. This suggests that in a spot labor market, there are no significant dynamic effects of the minimum wage increase, and the labor market adjusts immediately to the change. Finally, while Cengiz et al. (2019), which examines the U.S. labor market, finds that the estimated number of excess jobs and missing jobs are approximately equal, our analysis of the spot labor market reveals that the number of missing jobs ($-3.0\%$) exceeds that of excess jobs ($1.2\%$).

### v2

@@ Abstract
This paper provides causal evidence on the short-run effects of minimum wage hikes in a spot labor market, using confidential contract-level data from Timee, a Japanese job-matching platform. Leveraging high-frequency variation and a bin-level difference-in-differences design, we find a 2 percentage point decline in employment within affected wage bins, with losses concentrated below the new threshold. Transportation reimbursement—a flexible nonwage amenity—remains largely unchanged, indicating limited pass-through. These findings suggest that spot labor markets adjust rapidly through wage compliance, with the cost burden borne by establishments rather than offset through reductions in job amenities.
@@ 1 Introduction
In this paper, we use confidential, contract-level data from Timee, a Japanese spot labor market platform, to provide causal evidence on the effects of minimum wage hikes on short-term employment and transportation reimbursements. The former captures the immediate response of firms to minimum wage increases, while the latter, as an easily adjustable nonwage amenity, enables us to examine potential pass-through effects of minimum wage changes on broader compensation structures beyond base wages. Understanding these two outcomes is essential for capturing the impact of minimum wage hikes, as they reflect both the direct and indirect margins through which firms adjust labor conditions in response to wage floor policies.
@@
Our findings reveal that the impact of the minimum wage increase on employment in the lower end of the wage distribution. We find that the increase resulted in approximately 2 percentage point decline in total employment among wage bins directly affected by the policy, with a 3 percentage point reduction in jobs below the new minimum wage and a 1 percentage point increase just above it. Consistently, the employment elasticity with respect to minimum wages computed from Cengiz et al. (2019) is $-0.387$. These effects materialize immediately after the policy takes effect and remain stable throughout the post-reform period. Unlike evidence from broader labor markets where excess jobs offset missing jobs (Cengiz et al. 2019), our setting shows that missing jobs consistently exceed excess jobs, resulting in a net negative employment effect. These results support the view that the spot labor market adjusts quickly to minimum wage hikes, primarily through employment reallocation.
@@
Finally, we use our approach to estimate the effect of minimum wage increases on specific occupational subgroups. By leveraging high-frequency vacancy data across diverse job types, we detect meaningful variation in how employment adjusts within the wage distribution across occupations. The most noticeable reductions in low-wage employment are concentrated in sectors such as restaurants, retail, and event staffing, while other occupations—including logistics, customer service, and light work—exhibit more balanced patterns with offsetting gains. Importantly, in office work, we find evidence of upward shifts in job postings, suggesting that employers may be attempting to upgrade their workforce quality by attracting more capable workers — what appears to be a “quality jump” in response to rising labor costs. These patterns highlight how even modest policy changes can lead to occupational reallocation and qualitative shifts in job structure.
@@ 1.1 Related Literature
The rest of the paper is structured as follows: Section 2 outlines the context of the spot labor market and the Japanese Minimum Wage Act with related studies. Section 3 introduces our data and illustrates motivating data patterns. Section 4 presents our identification strategy. Section 5 reports our empirical findings. Section 6 explores the heterogeneous effects across prefectures and occupations. Section 7 concludes the paper.
@@ 2.2 Spot work platform
We use a term “spot work platform” to refer non-regular labor market provided by Timee, the huge online platform in Japan. [^7] According to the Japan Spot Work Association (JASWA), “spot work” refers to employment contracts for short-hour, one-off work engagements that are facilitated through digital platforms. These platforms serve as intermediaries that match job offers with available workers—often on a same-day or near-term basis—and rely on digital technologies to streamline the processes of matching, contracting, and compensation. JASWA plays a central role in promoting sound practices in the sector by issuing guidelines for legal compliance, offering interpretations of labor laws, and supporting proper labor management. In particular, it emphasizes the use of digital infrastructure to reduce the administrative burden traditionally associated with employment, including time tracking, wage calculation, and performance evaluation. This description draws on information provided by the association’s official overview and published guidelines (https://www.jaswa.or.jp/about/; https://jaswa.or.jp/pdf/spotwork_gudeline_v2.pdf). This platform operates as an on-demand staffing platform designed to connect businesses with temporary workers for short-term jobs through the app. The primary users of the platform are individuals seeking flexible, task-based employment rather than long-term positions and establishments seeking for short-term workers. Workers can register on the platform for free and immediately gain access to job offers from various companies, with the platform streamlining the process of matching workers with available shifts across industries such as food, retail, and logistics.
@@ 4.1 Prefecture-month-wage-bin level DID estimate
$\frac{Y_{j(p),t}}{N_{p,t}}=\sum_{l\neq-1}\sum_{e\neq\infty}\mu_{l,e}\mathbf{1}\{t-7=l\}\mathbf{1}\{E_{j(p)}=e\}+\alpha_{e}+\lambda_{t}+\epsilon_{j,t},$ (1)
@@
$\mu_{l,e}=E\biggr{[}\frac{Y_{j(p),7+l}}{N_{p,7+l}}-\frac{Y_{j(p),6}}{N_{p,6}}|E_{j(p)}=e\biggr{]}-E\biggr{[}\frac{Y_{j(p),7+l}}{N_{p,7+l}}-\frac{Y_{j(p),6}}{N_{p,6}}|E_{j(p)}=\infty\biggr{]},$
@@ 4.2 Excess-jobs-Missing-jobs Approach
Next, following Cengiz et al. (2019), we estimate the effect of minimum wage increase on the outcome throughout the wage distribution by aggregating each $\mu_{l,e}$, given by the prefecture-month-wage-bin level DID estimates. Let $\Delta a_{l}=\sum_{e=0}^{3}\mu_{l,e}$ and $\Delta b_{l}=\mu_{l,-1}$ denote the excess jobs and the missing jobs at a give relative period $l$, respectively. Let $\Delta a=\frac{1}{6}\sum_{l=0}^{5}\Delta a_{l}$ and $\Delta b=\frac{1}{6}\sum_{l=0}^{5}\Delta b_{l}$ denote the total excess jobs and missing jobs after the increase of minimum wage, respectively. Also, let $\Delta a_{e}=\frac{1}{6}\sum_{l=0}^{5}\mu_{e,l}$ and $\Delta b_{-1}=\frac{1}{6}\sum_{l=0}^{5}\mu_{-1,l}$ define the excess jobs and the missing jobs at a give relative exclusive group $e$, respectively. We compute the overall employment effect at a give relative period $l$, denoted as $\Delta e_{l}$, by summing the excess and missing jobs at that period: $\Delta e_{l}=\Delta a_{l}+\Delta b_{l}$. Furthermore, we calculate the total employment effect after the minimum wage increase, denoted as $\Delta e$, by summing $\Delta e_{\l}$ over all relevant periods: $\Delta e=\sum_{l=0}^{5}e_{l}$.
@@ 5.1 Prefecture-month-wage-bin Level DID Estimate on Employment
Three key findings emerge from Panel (a) and Panel (b). First, immediately after the minimum wage increase, there is a clear occurrence of job losses (missing jobs) below the minimum wage and job creation (excess jobs) above the minimum wage in the wage distribution. Second, the estimated excess jobs and missing jobs remain stable across months following the minimum wage increase. This suggests that in a spot labor market, there are no significant dynamic effects of the minimum wage increase, and the labor market adjusts immediately to the change. Finally, while Cengiz et al. (2019), which examines the U.S. labor market, finds that the estimated number of excess jobs and missing jobs are approximately equal, our analysis of the spot labor market reveals that the number of missing jobs ($-3.0\%$) exceeds that of excess jobs ($1.2\%$). [^13] The elasticity with respect to minimum wages computed from the formula of Cengiz et al. (2019) is $-0.387(0.132)$ where the bracket number is the standard error. See the detail in Appendix A.
@@ 5.3 Platform Revenue
6 Heterogeneous Job-Posting Responses to the Minimum Wage
6.1 Prefecture-month and Occupation-month level DID estimate
In section 5.1, we estimate the effect of the minimum wage increase on the outcome by aggregating the events of the minimum wage increase across prefectures. However, the minimum wage bite varies across prefectures, and the relationship between the employment effect and the Kaitz index (the ratio between the median wage and the minimum wage) is intrinsically interesting. We implement the DID identification strategy separately for each prefecture $p$, and show that the estimated employment effects are more negative in prefectures with a higher Kaitz index.
Specifically, we first construct a separate data set for each prefecture $p$ and run the following regression within each data set:
$Y^{p}_{j,t}=\sum_{l\neq-1}\sum_{e\neq\infty}\mu^{p}_{l,e}\mathbf{1}\{t-7=l\}\mathbf{1}\{E^{p}_{j}=e\}+\alpha^{p}_{e}+\lambda^{p}_{t}+\epsilon^{p}_{j,t}.$ (2)
Here, the superscript $p$ indicates that this regression is estimated separately for each prefecture $p$. As in (1), each coefficient $\mu^{p}_{l,e}$ is the DID estimand, capturing the dynamic effect of the minimum wage increase on the outcome at a given period $l$ in prefecture $p$ under the DID identifying assumptions:
$\mu^{p}_{l,e}=E[Y^{p}_{j,7+l}-Y^{p}_{j,6}|E^{p}_{j}=e]-E[Y^{p}_{j,7+l}-Y^{p}_{j,6}|E^{p}_{j}=\infty],$
where the pre-treatment period is $l=-1$ and the control group is group $\infty$ in prefecture $p$.
Next, as in Section 4.2, we compute the missing jobs, excess jobs, and the overall employment effect for each prefecture $p$ by aggregating $\mu^{p}_{l,e}$ in the fist step:
$\Delta a^{p}=\sum_{l=0}^{5}\Delta a^{p}_{l},\hskip 5.69054pt\Delta b^{p}=\sum_{l=0}^{5}\Delta b^{p}_{l},\hskip 5.69054pt\Delta e^{p}=\sum_{l=0}^{5}e^{p}_{l},$
where the superscript $p$ indicates that each quantity is defined within prefecture $p$.
Prefecture-month level
Figure 11 illustrates the relationship between the Kaitz index and employment outcomes across prefectures. The Kaitz index is calculated for each prefecture as the ratio of the October 2023 median wage to the increased minimum wage. Panel (a) plots the estimated numbers of missing and excess jobs against the Kaitz index. It shows a positive association between the Kaitz index and excess jobs, and a negative association with missing jobs. When the minimum wage is high relative to the median wage, minimum wage hikes tend to have a stronger impact. Panel (b) presents the relationship between the Kaitz index and the overall employment effect, revealing a negative correlation. These findings suggest that, in Japan’s spot work market, minimum wage increases have already begun to produce modest negative employment effects, in contrast to the United States, where such effects remain limited (Cengiz et al. 2019).
Figure 11: Relationship between Excess Jobs, Missing Jobs, Employment Change, and the Kaitz Index (Minimum-to-Median Wage Ratio) across Prefectures
Note: Figure 11 shows the binned scatter plots for missing and excess jobs in Panel (a), and total employment changes in Panel (b) by the value of the minimum-to-median wage ratio (Kaitz index) for the 47 prefecture-specific estimates. The Kaitz index is calculated for each prefecture as the ratio of the October 2023 median wage to the increased minimum wage.
Occupation-month level
While prior studies have often concentrated on sectors like restaurants—where the minimum wage is most binding and effects are more easily detected—we adopt a distributional approach that captures wage and employment responses even in industries with limited direct exposure. This enables a more comprehensive evaluation of the policy’s impact across a broader range of occupations. [^14] Lordan and Neumark (2018) and Aaronson and Phelan (2019) find that the minimum wage reduces the employment of the lowest-wage workers in routine automatable occupations. Dustmann et al. (2022) find that minimum wages could also affect productivity by improving the composition of jobs in the economy and in response to minimum wage increases, small, inefficient firms exited the market, and workers at these firms found employment at more productive firms. Also, Forsythe (2023) finds that there is no heterogeneity across occupation.
Figure 12 illustrates heterogeneous occupation-level responses to minimum wage hikes, revealing distinct patterns across job categories. The most pronounced employment losses below the new minimum wage occur in Restaurant, Retail, and Event Staff, where a substantial number of low-wage jobs disappear. In contrast, occupations such as Logistics, Customer Service, and Light Work exhibit relatively balanced patterns, with moderate excess job creation at or just above the new minimum that partially offsets missing jobs. Professional and Entertainment roles show smaller overall impacts, suggesting limited exposure to the minimum wage threshold. Notably, Office Work stands out as the only occupation with significant employment gains concentrated in the higher wage bins, implying that employers may be attempting to upgrade their workforce quality by attracting more capable workers—what appears to be a “quality jump” in response to rising labor costs. Overall, these findings suggest that minimum wage increases lead not only to occupational reallocation but also to qualitative adjustments in hiring, especially in tasks related to office work.
Figure 12: Impact of Minimum Wages on Employment and Wages by Occupations
Note: Figure 12 shows the estimated impact of the minimum wage increase on employment for each group $e$ in the wage distribution for each occupation in the platform.
@@ 7 Conclusion
Appendix A Appendix
A.1 Elasticities Related to the Minimum Wages
Following the method of Cengiz et al. (2019), we estimate (i) the elasticity of employment with respect to minimum wages and (ii) the employment elasticity with respect to the affected wages.
Table 2: Impact of Minimum Wages on Employment and Wages
Estimate (SE)
Missing jobs below new MW -0.03
Excess jobs above new MW 0.012
Affected wages 0.366
Affected employment -0.268
Employment elasticity w.r.t MW -0.387
Emp. elasticity w.r.t affected wage -0.732
Job below new MW 0.068
% MW changes 0.047
Note: The table reports the effects of a minimum wage increase based on the event study analysis exploiting 47 prefecture-level minimum wage changes in 2023 (i.e., about 4% minimum wage increase). The table reports six-months averaged post-treatment estimates on missing jobs up to 100JPY below the new minimum wage, excess jobs at and up to 300JPY above it, employment, and wages. Robust standard errors in parentheses are clustered by wage-bin separated by 10 JPY.
Table 2 reports six-month–averaged post-treatment effects for the baseline specification depicted in the Figure 7. The magnitude of missing jobs below new minimum wages exceeds the magnitude of excess jobs above new minimum wages in the six-month window. The estimated employment elasticity with respect to the minimum wage is -$0.387$, which is consistent with the Figure 7. Unlike the findings of Cengiz et al. (2019) and Dube and Zipperer (2024), our own wage elasticity is -$0.732$, indicating that wages are relatively elastic to wage adjustment related to the minimum wages compared by the median one (-$0.13$) shown by Dube and Lindner (2024).